Teresa deposits $4,000 in an account that earns 7% simple interest paid annually. How long will it be before the total amount is $6,500. Round your answer to the nearest year
Answer by CubeyThePenguin(3113)

A = P(1 + rt)

A = final amount = $6,500
P = principal (initial) amount = $4000
r = interest rate = 0.07
t = time (in years) = x

6500 = 4000(1 + 0.07x)
1.625 = 1 + 0.07x
0.625 = 0.07x
x = 8.928 years ----> 9 yrs